OpenStudy (anonymous):

Which of the following are rational numbers? A; √3/27 B; 3-27 (I already know this one is rational) C; √3 - √27 D; √3/ √27 E; ( √27+3)( √27-3)

OpenStudy (anonymous):

some gimmick here \[\frac{\sqrt{3}}{\sqrt{27}}=\sqrt{\frac{3}{27}}=\sqrt{\frac{1}{9}}\]

OpenStudy (anonymous):

which as you can see is \(\frac{1}{3}\) so pick that one for sure

OpenStudy (anonymous):

try multiplying out \[(\sqrt{27}+3)(\sqrt{27}-3)\] and see what you get

OpenStudy (anonymous):

So √3/√27 = 1/9 ?

OpenStudy (anonymous):

no it is \(\frac{1}{3}\)

OpenStudy (anonymous):

since the square root of 9 is 3

OpenStudy (anonymous):

oh ok. and (√27+3)(√27-3) = 39.5884572681199

OpenStudy (anonymous):

it is \[27-9=18\]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\[(a+b)(a-b)=a^2-b^2\]

OpenStudy (anonymous):

and so \[(\sqrt{27}+3)(\sqrt{27}-3)=27-9\]

OpenStudy (anonymous):

ok i kind of understand. is that all?

OpenStudy (anonymous):

yeah i think that is all the rationals ones